OpenStudy (anonymous):

If sin(theta) =4/5 and is in quadrant 2, the value of cot(theta)....Help Dx

OpenStudy (anonymous):

if sin(theta)= 4/5, then tan(theta)=4/3, but, since it's in quadrant 2, cos(theta)=-4/3. Then, cot(theta)=1/tan(theta) cot(theta)=-3/4

OpenStudy (anonymous):

sorry, I mean tan(theta) not cos(theta)

OpenStudy (anonymous):

Thank you :)